OpenStudy (anonymous):

15/36 in the lowest from

OpenStudy (fibonaccichick666):

well, do 15 and 36 have any common factors? Can you divide them by the same number?

OpenStudy (anonymous):

15/36=(3*5)/(3*12)=5/12

OpenStudy (anonymous):

I don't think so @FibonacciChick666

OpenStudy (fibonaccichick666):

do me a favor, write down the factors for both

OpenStudy (fibonaccichick666):

so what numbers can you multiply to get 15?

OpenStudy (anonymous):

3*5, 15*1

OpenStudy (fibonaccichick666):

goo, now how about 36?

OpenStudy (anonymous):

36*1 ,6*6, 12*3

OpenStudy (fibonaccichick666):

ok, now are there any numbers they both have?

OpenStudy (anonymous):

3 and 1

OpenStudy (fibonaccichick666):

good, so 1 doesn't help us much but 3 will,

OpenStudy (anonymous):

okay so now what

OpenStudy (fibonaccichick666):

In order to reduce the fraction, we have to "multiply by 1" (but it should be an improper fraction. So in this case we want to multiply by \[\frac{\frac{1}{3}}{\frac{1}{3}}\]

OpenStudy (fibonaccichick666):

ie: divide the top and the bottom by 3

OpenStudy (fibonaccichick666):

So, what does that change \(\frac{15}{36}\) into?